# Integrated Sensing and Communications Over the Years: An Evolution Perspective

### v2

@@ IV-C ISAC Datasets
SDP[200] • CSI
• Kinect data
Time >400 hours IEEE ISAC-ETI
@@
Radar Signatures of Human Activities[202] • Radar $\times$ $\times$ $\times$ $\times$ - University of Glasgow
@@
MmWave Gesture Dataset[205] • mmWave
@@
• Depth maps $\times$ $\times$ $\times$ $\times$
Time >1357 minutes Beijing university of Post and Telecommunications
DeepSense 6G [206] • RGB
@@
IV-C1 High-Quality Datasets
  • SDP[200]: Launched in February 2024 by IEEE ISAC-ETI, SDP is the industry’s first large-scale field-measured sensing dataset. It addresses major wireless sensing challenges by publicly releasing RF datasets and promoting high-performance and reliable wireless sensing solutions. It covers seven application scenarios, spans over 400 hours and 14.1 TB of data, and supports tasks including detection (e.g., presence of people or pets), positioning and tracking (humans, vehicles, drones), recognition (gesture and action), vital sign monitoring (breathing, heartbeat, sleep), and imaging (skeleton and environment reconstruction).
  • WiMANS[201]: A Wi-Fi-based multi-user activity sensing benchmark, WiMANS provides 9.4 hours of dual-band CSI synchronized with RGB videos of up to five users performing nine daily activities. It targets multi-user activity recognition, tracking, and pose estimation, addressing the limitations of single-user datasets.
  • Radar signatures of human activities[202]: This dataset, provided by the University of Glasgow, focuses on radar-based human activity recognition. It includes a range of human activity radar signals collected using FMCW radar, supporting the development and evaluation of radar signal processing and machine learning algorithms for human activity classification.
  • OPERAnet[204]: OPERAnet is a multi-modal dataset designed for indoor Human Activity Recognition (HAR) and localization. The dataset contains approximately 8 hours of data collected across two indoor environments with six participants performing six daily activities. It is designed to facilitate multi-modal research in smart homes, elderly care, and surveillance, with high-resolution activity and location labels.
  • MmWave Gesture Dataset[205]: MmWave-gesture-dataset features over 56,420 tracking samples and 1,357 minutes of gesture data, collected from two scenarios with varying distance ranges. It includes both preset and unexpected gestures, offering multi-dimensional data from raw signals to advanced processing information, which aids in gesture recognition technology development.
    Figure 11: Communication Security and Sensing Privacy in ISAC Systems
  • DeepSense 6G[206]: Designed to support research in telecommunications and sensory technologies for 6G networks, DeepSense 6G dataset provides a rich collection of real-world, multi-modal data. It captures communication and sensory inputs across diverse scenarios and times of day, facilitating deep insights into 6G technology performance and adaptability.
  • WALDO[207]: Representing a significant advancement in indoor positioning research, the WALDO dataset combines IEEE 802.11ay PHY software with multi-static radar systems. It provides radar and wireless signal data for developing advanced positioning algorithms capable of accurately estimating human positions in complex indoor environments.
@@ IV-C2 Summary and Discussion
Simulation datasets provide controllable environments for analyzing system performance and optimizing algorithm design. However, they often fail to capture the complexity and dynamics of real-world environments. By contrast, field-measured datasets such as SDP and DeepSense 6G, though resource-intensive to construct, provide higher fidelity, diversity, and representativeness. Such datasets are indispensable for training AI models that generalize well to practical ISAC scenarios, thereby accelerating progress in ISAC and edge intelligence technologies.
@@ IV-D2 How Can Massive, Heterogeneous Sensing Data Be Integrated at Scale
The fusion and processing of heterogeneous, multi-dimensional sensing data demand new algorithmic frameworks that can adapt to device diversity and coupled application requirements. Within a single coverage area, sensing data are inherently interdependent, necessitating efficient mechanisms for distributed model parameter generation. However, current distributed learning approaches often overlook device heterogeneity, resulting in inefficiencies and degraded performance. In massive machine-type communication (mMTC) scenarios, resource-constrained sensors may lack the capacity to perform local training, which forces trade-offs between S&C under stringent resource budgets [208]. Developing scalable, adaptive, and resource-aware data fusion strategies remains an open and critical challenge.
@@ V ISAC Security and Privacy
To fully explore the potential of ISAC technology in next-generation networks, the data security and sensing privacy should be well protected as illustrated in Fig. 11. The integration of S&C functions introduces new vulnerabilities due to dual-domain exposure. While sensing enhances communication, it also inadvertently exposes sensitive information, such as the location and movement patterns of targets. Furthermore, the shared infrastructure supporting both S&C increases the risk of data interception, heightening the likelihood of security breaches. Nonetheless, this integration also opens opportunities for designing enhanced security and privacy mechanisms.
@@ V-A Communication Security
Physical layer security (PLS) remains vital in ISAC systems, even if communication data is encrypted at higher layers. PLS offers additional protection against low-level threats and strengthens overall system security by preventing attacks that could compromise communication channel data. Conventional methods like secure beamforming are insufficient as they compromise sensing functionality. This necessitates the development of novel, tailored security approaches.
@@ V-A1 Artificial Noise (AN) Design
AN enhanced PLS by injecting interference to degrade the SINR at eavesdroppers, thus reducing their decoding ability while preserving communication and sensing quality for legitimate users. In [209], an optimization framework for AN-aided secure transmit beamforming in multi-user MISO ISAC systems maximized the secrecy rate by jointly designing beamformers and AN. It characterized the trade-offs between secrecy rate, communication quality, and sensing functionality and evaluated robustness against imperfect S&C CSI, confirming the effectiveness of AN in mitigating communication security threats.
@@ V-A2 Directional Modulation (DM)
DM uses symbol-level precoding to apply constructive interference (CI) to legitimate receivers while limiting signals at potential eavesdroppers by employing destructive interference (DI). [210] applied the CI-DI approach to mmWave frequencies, where angular correlations between S&C channels are critical. This technique effectively ensures data security by limiting interference from eavesdroppers, particularly in high-frequency bands.
@@ V-B Sensing Privacy
The integration of sensing functionality into wireless systems introduces additional privacy risks. Sensing signals emitted by dual-function BSs are reflected by targets. It may be intercepted by unauthorized receivers and exposing sensitive target-related information and compromising privacy. Unlike communication signals, which benefit from higher-layer encryption, passive sensing signals cannot be effectively encrypted, presenting a unique security challenge. Additional concerns arise in imaging and environmental sensing, where unencrypted control signals (e.g., pilot signal) may be exploited to infer user positions or construct environmental maps. These vulnerabilities underscore the need for robust privacy-preserving mechanisms.
@@ V-B1 Beamforming for Privacy
Innovative beamforming techniques are essential for preventing unauthorized data extraction while preserving S&C functionality. In [211], a beamforming-based approach was introduced to secure sensing functionality by incorporating AN. The method optimizes the beamforming and AN covariance matrix to enhance MI for legitimate radar users while reducing it for unauthorized entities. This effectively limits the adversary’s ability to intercept sensing data, protecting environmental privacy without compromising sensing accuracy for trusted users. Additionally, [212] proposed a privacy-preserving countermeasure using AI-supported signal obfuscation via adaptive beamforming. This method suppresses unauthorized sensing while maintaining the effectiveness of legitimate sensing operations.
@@ V-B2 CSI-Based User Location Anonymization
This method focuses on mitigating unauthorized positioning or tracking by obfuscating CSI changes. The goal is to obscure the user’s location while preserving sensor accuracy. In [213], an AI-based signal generation method is proposed to prevent unauthorized devices from extracting CSI-based location data, resulting in a 70% reduction in the risk of inference. Similarly, [214] proposed a multi-antenna signal-masking technique that exploits CSI time correlation and controls phase differences to thwart eavesdropping. Additionally, [215] introduced a pilot shielding method with tailored spectral characteristics to conceal genuine CSI variations, misleading eavesdroppers, and safeguarding location privacy.
@@ V-C Sensing-Assisted Communication Security
Embedding sensing functionality also provides unique opportunities to enhance PLS for communication. Leveraging its sensing capabilities, ISAC systems can dynamically estimate the sensing channel, i.e., the wiretap channel, by determining the angle of arrival of each target. This enables the system to prevent the target from decoding confidential information, thereby overcoming the traditional difficulty of obtaining eavesdropping CSI.
@@ V-C1 Sensing-Assisted PLS
By dynamically estimating the wiretap channel, ISAC systems can precisely track eavesdroppers’ locations and adapt communication strategies accordingly. This allows for real-time estimation of eavesdropper angles, enabling secure beamforming strategies that degrade the signal at the eavesdropper’s location while boosting it for legitimate users. Such methods, as proposed in [216] and [217], explore various optimization frameworks that balance communication performance with security objectives, demonstrating the mutual benefits of S&C functionalities in enhancing PLS. In [218], the iterative optimization of beamforming and AN strategies results in improved secrecy rates by minimizing the potential for eavesdropping while dynamically adjusting to the eavesdropper’s estimated location.
@@ V-C2 Sensing-Assisted Covert Communication
Unlike defending information eavesdropping in PLS, covert communication aims to prevent adversaries from detecting the existence of communication itself, requiring high-level confidentiality protection. Sensing-assisted covert communication leverages the dual functionality of radar and communication systems to ensure that information transmission remains undetectable by adversaries. These systems use sensing to track eavesdroppers and prevent communication signal detection by optimizing transmission characteristics, such as beamforming and power allocation. One approach involves jammer-assisted techniques, where a jammer is deployed to obscure transmission while the radar detects targets and communicates covertly. Optimization methods such as SDR [219] maximize covert throughput while maintaining radar detection constraints. Joint radar and communication beamforming designs also optimize covert throughput while maintaining radar performance. However, using dual-function waveforms introduces trade-offs between radar detection and covert transmission, balancing communication secrecy with efficient target detection.
Figure 12: Brief Review of 3GPP Standards Progress in ISAC
@@ V-D1 How to Model and Defend Against Hybrid Cross-Domain Attacks
V-D2 How to Validate ISAC Security Frameworks under Real-World Constraints
Most existing studies have relied on mathematically intensive optimization strategies, typically assuming perfect knowledge of attacker locations, CSI, and environmental conditions. However, real-world ISAC systems will operate under imperfect CSI, dynamic environments, and hardware limitations. For instance, [221] explored UAV-assisted ISAC security. However, practical challenges (such as power, latency, and actual sensing precision) remain largely theoretical and lack empirical validation in practice.
@@ VI Standardization Progress from Third Generation Partnership Project (3GPP) to ITU
Standardization has played a pivotal role in shaping the ISAC landscape, from addressing interoperability challenges and enabling widespread adoption to accelerating technological innovation. This section reviews recent progress in ISAC standardization, with particular focus on the collaborative roles of 3GPP, IEEE, and ITU. Key contributions on NR air interface advancements, emerging technologies, and regulatory frameworks are also examined.
@@ VI-A Efforts and Trends in Standardization
The IEEE, alongside key organizations such as the ITU and 3GPP, played complementary roles in the development and standardization of NR air interface technologies.
@@ VI-A1 Brief Review of 3GPP Standards
Releases 15–17: Foundational Enablers
Although ISAC was not explicitly specified in its early releases, multiple enabling technologies were standardized in Rel 15–17. Rel 15 established the standalone 5G NR architecture with massive MIMO, beamforming, and flexible spectrum utilization [222, 223], providing the structural basis for joint operations. Release 16 advanced positioning via multi-round trip time (Multi-RTT), uplink (UL)-AoA, and downlink angle-of-departure (DL-AoD) [224], complemented by refined positioning reference signal (PRS) and sounding reference signal (SRS) mechanisms, as well as mmWave operation for cellular vehicle-to-everything scenarios. Release 17 further expanded the mmWave bandwidth up to 400 MHz, enabling sensing resolution at the sub-meter level, and specified industrial-grade positioning requirements of 20–30 cm [225]. In addition, advanced AoA/AoD estimation methods, such as multiple signal classification, were aligned with the stringent precision requirements of ISAC systems.
@@
Release 18 introduced 5G Advanced, featuring RAN intelligence and carrier-phase–based positioning, which significantly improved sensing accuracy. Key enhancements included sidelink positioning, enabling vehicles and devices to collaborate on localization in complex environments. The 3GPP TS 38.355 specification defined mechanisms for exchanging positioning information between UE and networks, thereby enhancing V2X localization performance [226, 227]. This release also addressed multi-device coordination and dynamic resource allocation, laying the groundwork for more advanced ISAC integration in subsequent releases.
@@
In early 2022, Release 19 SA1 initiated a feasibility study on ISAC services, identifying key use cases and requirements for commercial deployment across multiple industries. By June 2023, the 3GPP SA plenary organized a workshop for Release 19 SA2, where ISAC was formally proposed by several industrial stakeholders and included as a potential study item. The 3GPP TR 22.837 identified 32 representative ISAC use cases, grouped into five deployment scenarios: smart transportation, smart industry, smart home, smart city, and smart drones [228].
In August 2023, 3GPP introduced TS 22.137, Service Requirements for ISAC, during the SA1#103 meeting [229]. This specification defined eight KPIs for wireless sensing in 5G, including positioning accuracy, velocity estimation, and sensing resolution. Subsequently, in December 2023, the RAN#102 meeting approved a dedicated study item on ISAC channel modeling, which was further refined at RAN#103 in March 2024 [230]. The objective of this study is to establish a unified framework for object detection and tracking, thereby improving the reliability of ISAC across diverse scenarios such as UAVs, pedestrians, vehicles, and hazardous objects. Building upon these standardization efforts, ISAC air interface technologies are expected to be systematically investigated in Release 20 beginning in 2025, with final specifications anticipated in Release 21 as part of the 6G standardization cycle [231].
@@
Enhancements in 5G infrastructure and air interface designs are expected to significantly strengthen wireless channels, thereby supporting more accurate and fine-grained environmental sensing. A representative example is Huawei’s proposal to integrate the sensing function into the 5G-Advanced core network, which aims to enable functions such as sensing control, AI-based data processing, secure data handling, and monetization of sensing data [232]. At the same time, the design of ISAC service processes and protocol stacks remains an open challenge. Interoperability requires well-defined signaling procedures for sensing capability negotiation, demand triggering, and result reporting. Whether existing protocol stacks can be extended to accommodate these new requirements, or if entirely new signaling frameworks must be developed, remains an open research problem. Addressing these issues will be essential to ensure seamless integration of ISAC into commercial deployments.

### v3

@@ IV-C ISAC Datasets
IV-C1 High-Quality Datasets
  • SDP[sdp8]: Launched in February 2024 by IEEE ISAC-ETI, the SDP platform is the industry’s first large-scale field-measured sensing dataset. This platform addresses major wireless sensing challenges by publicly releasing RF datasets and promoting high-performance and reliable wireless sensing solutions. The platform has released a large-scale field-measured dataset that covers 7 typical sensing application scenarios. It spans 400 hours, contains 14.1T of data, and can be used for algorithm research and model training for various sensing tasks. These tasks include detection (presence detection of targets like people and pets), positioning & tracking (tracking of people, vehicles, and drones), recognition (gesture and action recognition), vital signs (breathing, heartbeat, sleep), and imaging (human skeleton and environmental reconstruction).
  • WiMANS[huang2025wimans]: A Wi-Fi-based multi-user activity sensing benchmark, WiMANS provides 9.4 hours of dual-band CSI synchronized with RGB videos of up to five users performing nine daily activities. It targets multi-user activity recognition, tracking, and pose estimation, addressing the limitations of single-user datasets.
  • Radar signatures of human activities[UnivGlasgow]: This dataset, provided by the University of Glasgow, focuses on radar-based human activity recognition. It includes a range of human activity radar signals collected using FMCW radar, supporting the development and evaluation of radar signal processing and machine learning algorithms for human activity classification.
  • OPERAnet[bocus2022operanet]: OPERAnet is a multi-modal dataset designed for indoor Human Activity Recognition (HAR) and localization. The dataset contains approximately 8 hours of data collected across two indoor environments with six participants performing six daily activities. It is designed to facilitate multi-modal research in smart homes, elderly care, and surveillance, with high-resolution activity and location labels.
  • MmWave Gesture Dataset[liu2021m]: MmWave-gesture-dataset features over 56,420 tracking samples and 1,357 minutes of gesture data, collected from two scenarios with varying distance ranges. It includes both preset and unexpected gestures, offering multi-dimensional data from raw signals to advanced processing information, which aids in gesture recognition technology development.
@@ IV-C1 High-Quality Datasets
    SDP[sdp8] • CSI
    • Kinect data Size >14.1 TB Time >400 hours IEEE ISAC-ETI
@@
    Radar Signatures of Human Activities[UnivGlasgow] • Radar $\times$ $\times$ $\times$ $\times$ - University of Glasgow
@@
    MmWave Gesture Dataset[liu2021m] • mmWave
@@
    • Depth maps $\times$ $\times$ $\times$ $\times$ Size >56,420 Time >1357 minutes Beijing university of Post and Telecommunications
    DeepSense 6G [deepsense6g] • RGB
@@
  • DeepSense 6G[deepsense6g]: Designed to support research in telecommunications and sensory technologies for 6G networks, DeepSense 6G dataset provides a rich collection of real-world, multi-modal data. It captures communication and sensory inputs across diverse scenarios and times of day, facilitating deep insights into 6G technology performance and adaptability.
  • WALDO[nist-waldo]: Representing a significant advancement in indoor positioning research, the WALDO dataset combines IEEE 802.11ay PHY software with multi-static radar systems. It provides radar and wireless signal data for developing advanced positioning algorithms capable of accurately estimating human positions in complex indoor environments.
@@ IV-C2 Summary and Discussion
    Simulation datasets provide controllable environments for analyzing system performance and optimizing algorithm design. However, they often fail to capture the complexity and dynamics of real-world environments. By contrast, field-measured datasets such as SDP and DeepSense 6G, though resource-intensive to construct, provide higher fidelity, diversity, and representativeness. Such datasets are indispensable for training AI models that generalize well to practical ISAC scenarios, thereby accelerating progress in ISAC and edge intelligence technologies[wang2025surveywifisensinggeneralizability].
@@ IV-D2 How Can Massive, Heterogeneous Sensing Data Be Integrated at Scale
    The fusion and processing of heterogeneous, multi-dimensional sensing data demand new algorithmic frameworks that can adapt to device diversity and coupled application requirements. Within a single coverage area, sensing data are inherently interdependent, necessitating efficient mechanisms for distributed model parameter generation. However, current distributed learning approaches often overlook device heterogeneity, resulting in inefficiencies and degraded performance. In massive machine-type communication (mMTC) scenarios, resource-constrained sensors may lack the capacity to perform local training, which forces trade-offs between S&C under stringent resource budgets [ullah2018information]. Developing scalable, adaptive, and resource-aware data fusion strategies remains an open and critical challenge.
    Figure 11: Communication Security and Sensing Privacy in ISAC Systems
@@ V ISAC Security and Privacy
    To fully explore the potential of ISAC technology in next-generation networks, the data security and sensing privacy should be properly addressed, as illustrated in Fig. 11. The integration of S&C functions introduces new vulnerabilities due to dual-domain exposure. While sensing enhances communication, it may also inadvertently exposes sensitive information, such as the location and movement patterns of targets. Furthermore, the shared infrastructure supporting both S&C increases the risk of data interception, thereby heightening the likelihood of security breaches. Nonetheless, this integration also opens opportunities for designing enhanced security and privacy mechanisms[yang2025privacy].
@@ V-A Communication Security
    Physical layer security (PLS) remains vital in ISAC systems, even if communication data is encrypted at higher layers. PLS offers additional protection against low-level threats and strengthens overall system security by preventing attacks that could compromise communication channel data. Traditional secure beamforming approaches become insufficient because they may compromise sensing functionality. This necessitates the development of novel, tailored security approaches.
@@ V-A1 Artificial Noise (AN) Design
    AN enhanced PLS by injecting interference to degrade the SINR at eavesdroppers, thus reducing their decoding ability while preserving communication and sensing quality for legitimate users. In [9199556], an optimization framework for AN-aided secure transmit beamforming in multi-user MISO ISAC systems maximized the secrecy rate by jointly designing beamformers and AN. It characterized the trade-offs between secrecy rate, communication quality, and sensing functionality and demonstrated robustness under imperfect S&C CSI, confirming the effectiveness of AN in mitigating communication security threats.
@@ V-A2 Directional Modulation (DM)
    DM uses symbol-level precoding to apply constructive interference (CI) to legitimate receivers while limiting signals at potential eavesdroppers by employing destructive interference (DI). [9737364] applied the CI-DI approach to mmWave frequencies, where angular correlations between S&C channels are critical. This technique effectively enhances PLS by restricting signal leakage toward unauthorized directions, reducing energy consumption and system cost in high-frequency ISAC deployments.
@@ V-B Sensing Privacy
    The integration of sensing functionality into wireless systems introduces additional privacy risks. Sensing signals emitted by dual-function BSs are reflected by targets. It may be intercepted by unauthorized receivers and exposing sensitive target-related information and compromising privacy. Unlike communication signals, which benefit from higher-layer encryption, passive sensing signals cannot be effectively encrypted, presenting a unique security challenge. Moreover, measures to mitigate interference between S&C functions can inadvertently reveal critical transmitter information, such as location. Enhanced sensing capabilities further elevate the risk of passive eavesdropping by unauthorized nodes. Additional concerns arise in imaging and environmental sensing, where unencrypted control signals (e.g., pilot signal) may be exploited to infer user positions or construct environmental maps. These vulnerabilities underscore the need for robust privacy-preserving mechanisms.
@@ V-B1 Beamforming for Privacy
    Innovative beamforming techniques are crucial for preventing unauthorized data extraction while maintaining S&C functionality. In [10587082], a beamforming-based method was proposed to secure sensing functionality by integrating AN. It optimized the beamforming and AN covariance matrix to enhance MI for legitimate radar users, while deteriorating it for unauthorized ones. This successfully restricts the adversary’s ability to intercept sensing data, protecting environmental privacy without degrading sensing accuracy. Furthermore, [10574259] proposed an AI-assisted adaptive beamforming method to obfuscate signals, blocking unauthorized sensing while keeping legitimate activities functional. This method suppresses unauthorized sensing while maintaining the effectiveness of legitimate sensing operations.
@@ V-B2 CSI-Based User Location Anonymization
    This method focuses on mitigating unauthorized positioning or tracking by obfuscating CSI changes. The goal is to obscure the user’s location while preserving sensor accuracy. In [Wang2024GenerativeAB], an AI-based signal generation method is proposed to prevent unauthorized devices from extracting CSI-based location data, resulting in a 70% reduction in the risk of inference. Similarly, [10556606] proposed a multi-antenna signal-masking technique that exploits CSI time correlation and controls phase differences to thwart eavesdropping. Additionally, [10611746] introduced a pilot-shielding method with tailored spectral characteristics to conceal genuine CSI variations, misleading eavesdroppers, and safeguarding location privacy.
@@ V-C Sensing-Assisted Communication Security
    While the integration of sensing functionality in ISAC systems introduces new challenges in ensuring communication security and protecting sensor privacy, it also provides unique opportunities to enhance PLS. Leveraging its sensing capabilities, ISAC systems can dynamically estimate the sensing channel, i.e., the wiretap channel, by determining the angle of arrival of each target. This enables the system to prevent the target from decoding confidential information, thereby overcoming the traditional difficulty of obtaining eavesdropping CSI, a well-established limitation in conventional communication-only systems.
@@ V-C1 Sensing-Assisted PLS
    Sensing-assisted PLS represents a promising approach for enhancing the security of ISAC systems, where the sensing functionality aids in improving the system’s ability to thwart eavesdropping and information leakage. By dynamically estimating the wiretap channel, ISAC systems can precisely track eavesdroppers’ locations and adapt communication strategies accordingly. This is particularly beneficial in overcoming the challenges posed by traditional methods that require difficult-to-obtain eavesdropping CSI. This allows for real-time estimation of eavesdropper angles, enabling secure beamforming strategies that degrade the signal at the eavesdropper’s location while boosting it for legitimate users. Such methods, as proposed in [10349846] and [xu2024sensingaidednearfieldsecurecommunications], explore various optimization frameworks that balance communication performance with security objectives, demonstrating the mutual benefits of S&C functionalities in enhancing PLS. In [9838753], the iterative optimization of beamforming and AN strategies results in improved secrecy rates by minimizing the potential for eavesdropping while dynamically adjusting to the eavesdropper’s estimated location. By combining S&C capabilities, these systems represent a significant advancement in securing the wireless transmission of sensitive information in next-generation ISAC networks.
@@ V-C2 Sensing-Assisted Covert Communication
    Unlike defending information eavesdropping in PLS, covert communication aims to prevent adversaries from detecting the existence of communication itself, requiring high-level confidentiality protection. Sensing-assisted covert communication leverages the dual functionality of radar and communication systems to ensure that information transmission remains undetectable by adversaries. These systems use sensing to track eavesdroppers and prevent communication signal detection by optimizing transmission characteristics, such as beamforming and power allocation. One approach involves jammer-assisted techniques, where a jammer is deployed to obscure transmission while the radar detects targets and communicates covertly. Optimization methods such as SDR [10507294] maximize covert throughput while maintaining radar detection constraints. Joint radar and communication beamforming designs also optimize covert throughput while maintaining radar performance. However, using dual-function waveforms introduces trade-offs between radar detection and covert transmission, balancing communication secrecy with efficient target detection. These approaches highlight ISAC’s potential to offer secure, covert communication by balancing the competing demands of S&C performance.
@@ V-D1 How to Model and Defend Against Hybrid Cross-Domain Attacks
    V-D2 How to Validate ISAC Security Frameworks under RealWorld Constraints
    Most existing studies have relied on mathematically intensive optimization strategies, typically assuming perfect knowledge of attacker locations, CSI, and environmental conditions. In reality, ISAC systems must operate with imperfect CSI, mobility, hardware impairments, latency constraints, and limited sensing precision. Although [10529955] investigated UAV-assisted ISAC security, its experimental validation remains limited, highlighting the need for prototypes, datasets, hardware testbeds, and standardized evaluation procedures.
    Figure 12: Brief Review of 3GPP Standards Progress in ISAC
@@ VI Standardization Progress from Third Generation Partnership Project (3GPP) to ITU
    Standardization has played a pivotal role in shaping the ISAC landscape, addressing interoperability challenges, enabling widespread adoption, and accelerating technological innovation.This section reviews recent progress in ISAC standardization, with particular focus on the collaborative roles of 3GPP, IEEE, and ITU. Key contributions on NR air interface advancements, emerging technologies, and regulatory frameworks are also examined.
@@ VI-A Efforts and Trends in Standardization
    The IEEE, alongside key organizations such as the ITU and 3GPP, has played complementary roles in the development and standardization of NR air interface technologies.
@@ VI-A1 Brief Review of 3GPP Standards
    Release 15: Foundational NR Technologies
    Rel-15 marked the launch of the first complete standalone NR architecture. Foundational technologies, including beamforming, massive MIMO, and flexible spectrum utilization, were introduced. These advancements, outlined in standards such as TS 38.300 and TS 38.104 [TS38300, TS38104], laid the groundwork for ISAC by enabling efficient resource allocation and high-performance signal processing. Although ISAC was not explicitly addressed, these technologies provided critical building blocks for future integration.
    Release 16: Positioning and Multicast for ISAC
    Rel-16 brought improved system efficiency, expanded use cases, and additional spectrum for 5G. Though it did not directly address ISAC, it introduced key enablers. Notably, 3GPP TR 38.855 introduced advanced positioning methods, including multi-round trip time (Multi-RTT), uplink-AoA, and downlink angle-of-departure (DL-AoD))[TR38855], improving accuracy in both indoor and outdoor environments. The positioning reference signal (PRS) and sounding reference signal (SRS) refined time-of-arrival and uplink-based positioning, with Fig. 13 showing the Multi-RTT procedure. Rel-16 also added multicast capabilities for vehicular use, enabling C-V2X communication. Initial mmWave support in Rel-16 paved the way for ISAC’s evolution in future 5G and 6G systems.
    Release 17: mmWave ISAC Enhancements
    Rel-17 introduced advanced MIMO, improved coverage, spectrum extension, and satellite integration, forming the foundation for comprehensive 5G upgrades. In vertical industries like factory automation and transportation, 3GPP TS 22.104 imposed stringent positioning requirements, including 20–30 cm accuracy[TS22104]. Rel-17’s 5G mmWave systems achieved 400-MHz bandwidth, offering 0.1875-meter resolution[shi2022device]. Multi-antenna BSs exploit AoA/AoD estimation technologies such as multiple signal classification (MUSIC) for precise target detection. ISAC takes advantage of the extensive cellular infrastructure and high-frequency bands, reinforced by ultra-large-scale MIMO for increased bandwidth and superior angular resolution. By combining low-latency, high-precision communication with advanced sensing, ISAC facilitates remote object detection and environmental analysis without physical contact.
@@
    Rel-18 introduced 5G Advanced, featuring RAN intelligence and carrier-phase–based positioning, which significantly improved sensing accuracy. Key enhancements included sidelink positioning, enabling vehicles and devices to collaborate in localization in complex environments. The 3GPP TS 38.355 specification defined mechanisms for exchanging positioning information between UEs and networks, thereby enhancing V2X localization performance [3gpp-ts-38.355, 10517597]. This release also addressed multi-device coordination and dynamic resource allocation, laying the groundwork for more advanced ISAC integration in subsequent releases.
@@
    In early 2022, Rel-19 SA1 initiated a feasibility study on ISAC services, identifying use cases and requirements for commercial ISAC implementation across industries. By June 2023, the 3GPP SA plenary organized a workshop for Rel-19 SA2, where ISAC was proposed by several companies and included as a potential study item for Rel-19. TR 22.837 [3GPP-rel19] identified 32 ISAC use cases across five deployment scenarios: smart transportation, smart industry, smart home, smart city, and smart drones.
    In August 2023, 3GPP introduced TS 22.137, Service Requirements for ISAC, during the SA1#103 meeting [TS22137]. This specification defined eight KPIs for wireless sensing in 5G, including positioning accuracy, velocity estimation, and sensing resolution. Subsequently, in December 2023, the RAN#102 meeting approved a dedicated study item on ISAC channel modeling, which was further refined at RAN#103 in March 2024 [R1-240]. The objective of this study is to establish a unified framework for object detection and tracking, thereby improving the reliability of ISAC across diverse scenarios such as UAVs, pedestrians, vehicles, and hazardous objects. Building upon these standardization efforts, ISAC air interface technologies are expected to be systematically investigated in Rel-20 beginning in 2025, with final specifications anticipated in Rel-21 as part of the 6G standardization cycle [lin2023bridge6g5gadvancedevolution].
@@
    Enhancements in 5G infrastructure and air interface designs are expected to significantly strengthen wireless channels, thereby supporting more accurate and fine-grained environmental sensing. A representative example is Huawei’s proposal to integrate the sensing function into the 5G-Advanced core network, which aims to enable functions such as sensing control, AI-based data processing, secure data handling, and monetization of sensing data [Huawei2023]. At the same time, the design of ISAC service processes and protocol stacks remains an open challenge. Interoperability requires well-defined signaling procedures for sensing capability negotiation, demand triggering, and result reporting. Whether existing protocol stacks can be extended to accommodate these new requirements, or if entirely new signaling frameworks must be developed, remains an open research problem. Addressing these issues will be essential to ensure seamless integration of ISAC into commercial deployments.